OpenStudy (slickster256):

A sphere and a cylinder have the same radius and height. The volume of the cylinder is mc013-1.jpg. What is the volume of the sphere?

OpenStudy (danjs):

k

OpenStudy (danjs):

What is the volume of the cylinder?, it messed up

OpenStudy (slickster256):

50ft sorry it messed up

OpenStudy (danjs):

50 ft^3 ?

OpenStudy (slickster256):

yea

OpenStudy (danjs):

ok...

OpenStudy (danjs):

Volume of a Sphere = (4/3) * pi * r^3 and Volume Cyl = pi * r^2 * h

OpenStudy (danjs):

you recognize those?

OpenStudy (slickster256):

yes those are formulas for volume

OpenStudy (danjs):

They both have the same radius, r the thing to recognize is it says they are both the same height, so...

OpenStudy (danjs):

the height of the cylinder is the diameter of the sphere, or , 2*r

OpenStudy (danjs):

you see that?

OpenStudy (slickster256):

yes but r^2 and r^3 doeesnt simplify does it?

OpenStudy (danjs):

sub in h=2r for the cylinder

OpenStudy (slickster256):

oh....

OpenStudy (danjs):

Sphere Vol = (4/3) * pi * r^3 Cylinder Vol = 50 = pi * r^2 * 2r

OpenStudy (danjs):

Now you have 2 equations with 2 unknown values, r and sphere volume. you can solve for those

OpenStudy (slickster256):

thank you very much

OpenStudy (danjs):

2pi *r^3 = 50 , solve r, use that in sphere volume..

OpenStudy (danjs):

welcome

OpenStudy (danjs):

i got the volume sphere as 100/3 ft^3, for ya to check..